# TimeSeriesBench: An Industrial-Grade Benchmark for Time Series Anomaly Detection Models

### v1

@@ E.4. Performance, Cost, Efficiency Trade-off
Table 3. Performance comparison using naive learning schema under reduced-length $F1_{best}$ and $AUPRC$.
Method AIOPS NAB TODS WSD Yahoo UCR NEK
MatrixProfile 0.0281 0.0097 0.1769 0.0975 0.2058 0.1407 0.0371 0.0234 0.2051 0.1445 0.5066 0.4711 0.3163 0.2439
SAND 0.0674 0.0262 0.3036 0.2169 0.2544 0.1879 0.0952 0.0666 0.468 0.4343 0.5494 0.5344 0.1758 0.1304
Sub-LOF 0.3764 0.31 0.6982 0.5835 0.5791 0.5236 0.7194 0.6895 0.4875 0.4443 0.5806 0.5325 0.6807 0.6423
AR 0.7517 0.75 0.9035 0.8691 0.7472 0.7677 0.868 0.8609 0.9195 0.9057 0.4454 0.4026 0.9555 0.9647
LSTMAD-$\alpha$ 0.8081 0.8236 0.8984 0.8358 0.7015 0.7201 0.9381 0.9298 0.5496 0.5026 0.4836 0.441 0.9971 0.9959
LSTMAD-$\beta$ 0.8034 0.8219 0.9031 0.841 0.7589 0.7706 0.939 0.9309 0.5366 0.4868 0.4366 0.3931 0.9744 0.9716
AE 0.7222 0.7295 0.8751 0.8146 0.6842 0.6675 0.918 0.9049 0.612 0.5692 0.4198 0.3879 0.9801 0.9824
EncDecAD 0.7583 0.7621 0.8752 0.8136 0.6359 0.6234 0.9285 0.9214 0.5018 0.449 0.3178 0.2829 0.755 0.733
Donut 0.6957 0.6791 0.8397 0.7607 0.7207 0.7061 0.8787 0.8668 0.6813 0.6441 0.4409 0.3968 0.9885 0.9923
FCVAE 0.7851 0.8079 0.874 0.8539 0.7145 0.7162 0.9087 0.8966 0.6883 0.6498 0.5873 0.5482 0.8977 0.9128
SRCNN 0.1627 0.0972 0.5802 0.4946 0.3042 0.1931 0.1785 0.0993 0.1261 0.0662 0.3227 0.2667 0.4365 0.3501
AnomalyTransformer 0.3728 0.3053 0.7739 0.691 0.2827 0.182 0.2984 0.2128 0.145 0.0879 0.3104 0.2548 0.3997 0.2617
TFAD 0.3551 0.2529 0.5746 0.532 0.4526 0.3943 0.7561 0.7115 0.781 0.7501 0.3398 0.3077 0.625 0.5786
TranAD 0.6486 0.5798 0.9101 0.8627 0.2531 0.1988 0.6738 0.6089 0.562 0.516 0.2527 0.2199 0.9071 0.8928
TimesNet 0.6737 0.6674 0.8419 0.7845 0.3695 0.3327 0.8684 0.8466 0.4444 0.4089 0.2493 0.2142 0.9335 0.9388
OFA 0.6891 0.6973 0.8869 0.8468 0.6263 0.5683 0.8784 0.8683 0.7512 0.7145 0.4169 0.3804 0.9471 0.9553
FITS 0.7139 0.7187 0.8617 0.8168 0.5371 0.5119 0.8796 0.8822 0.7564 0.7235 0.4092 0.3657 0.8975 0.9032
Table 4. Performance comparison using all-in-one learning schema under reduced-length $F1_{best}$ and $AUPRC$.
Method AIOPS NAB TODS WSD Yahoo UCR NEK
MatrixProfile - - - - - - - - - - - - - -
SAND - - - - - - - - - - - - - -
Sub-LOF - - - - - - - - - - - - - -
AR 0.8214 0.8203 0.8881 0.8462 0.7859 0.8365 0.9276 0.9179 0.948 0.9392 0.4636 0.4182 0.9612 0.9719
LSTMAD-$\alpha$ 0.8141 0.8343 0.8492 0.7889 0.7538 0.7738 0.9467 0.9401 0.8 0.7829 0.6144 0.5598 0.9349 0.9528
LSTMAD-$\beta$ 0.8151 0.8377 0.8484 0.7821 0.7495 0.7771 0.9445 0.9391 0.8114 0.7922 0.6101 0.5602 0.961 0.9781
AE 0.7448 0.7523 0.8375 0.7793 0.7712 0.7441 0.9192 0.9138 0.8205 0.8018 0.3644 0.3208 0.9609 0.9757
EncDecAD 0.7003 0.6894 0.8336 0.7786 0.6534 0.6592 0.8353 0.8182 0.5842 0.5378 0.3333 0.2867 0.9534 0.9538
Donut 0.5827 0.5578 0.7966 0.7085 0.836 0.8381 0.7666 0.7271 0.6498 0.6161 0.4 0.3395 0.5801 0.4923
FCVAE 0.7593 0.7521 0.8857 0.8587 0.8526 0.8535 0.8121 0.7814 0.8537 0.8292 0.4766 0.418 0.8148 0.793
SRCNN 0.6672 0.632 0.635 0.5648 0.2598 0.1658 0.7742 0.7337 0.1728 0.1056 0.3791 0.3176 0.6173 0.5906
AnomalyTransformer 0.3656 0.2103 0.7322 0.6432 0.2631 0.1561 0.5838 0.4227 0.2377 0.1468 0.2372 0.1777 0.666 0.5539
TFAD 0.4889 0.4062 0.4679 0.3996 0.5626 0.5411 0.8016 0.7602 0.9361 0.9236 0.4941 0.4646 0.4103 0.2853
TranAD 0.6561 0.6049 0.8567 0.7921 0.5624 0.5291 0.7409 0.6995 0.5211 0.4643 0.2483 0.2181 0.9571 0.9524
TimesNet 0.4988 0.4657 0.8544 0.8055 0.5032 0.4293 0.539 0.4602 0.5341 0.5098 0.1856 0.149 0.9257 0.9459
OFA 0.5544 0.4993 0.861 0.7911 0.6187 0.6129 0.8111 0.7877 0.766 0.7414 0.2888 0.2334 0.7642 0.7389
FITS 0.6986 0.6966 0.8683 0.8196 0.7146 0.7446 0.8788 0.8745 0.9261 0.9189 0.428 0.391 0.9625 0.9719
Table 5. Performance comparison using zero-shot learning schema under reduced-length $F1_{best}$ and $AUPRC$.
Method AIOPS NAB TODS WSD Yahoo UCR NEK
MatrixProfile - - - - - - - - - - - - - -
SAND - - - - - - - - - - - - - -
Sub-LOF - - - - - - - - - - - - - -
AR 0.8013 0.7928 0.8967 0.8306 0.7975 0.8589 0.926 0.9066 0.9366 0.9234 0.487 0.443 0.9641 0.9723
LSTMAD-$\alpha$ 0.8184 0.8304 0.8593 0.7913 0.677 0.6834 0.9349 0.92 0.8277 0.8141 0.5927 0.539 0.9539 0.9692
LSTMAD-$\beta$ 0.8073 0.8233 0.8556 0.7915 0.593 0.5935 0.9357 0.9219 0.7942 0.7745 0.5603 0.5128 0.9613 0.9718
AE 0.7689 0.7787 0.8538 0.7917 0.7105 0.6926 0.8969 0.8841 0.8595 0.8437 0.3239 0.284 0.9591 0.9727
EncDecAD 0.7537 0.7694 0.85 0.7796 0.7089 0.6829 0.8306 0.8163 0.5958 0.5499 0.3209 0.2764 0.9441 0.9526
Donut 0.6863 0.6724 0.7187 0.603 0.6223 0.5933 0.7934 0.7568 0.7034 0.6733 0.4148 0.367 0.6204 0.5069
FCVAE 0.7983 0.8143 0.9385 0.8895 0.7987 0.7884 0.7707 0.7316 0.796 0.7683 0.4101 0.3642 0.7979 0.8044
SRCNN 0.6888 0.6507 0.6421 0.6205 0.2977 0.18 0.7414 0.6936 0.1437 0.0859 0.3582 0.2889 0.5261 0.4788
AnomalyTransformer 0.4361 0.2757 0.5601 0.4912 0.2199 0.1531 0.3027 0.2065 0.4302 0.3613 0.2749 0.2116 0.5376 0.4876
TFAD 0.4659 0.3392 0.2806 0.2419 0.6352 0.619 0.8902 0.8634 0.9312 0.916 0.4689 0.44 0.4527 0.3534
TranAD 0.4662 0.3843 0.9333 0.8698 0.4013 0.3717 0.62 0.5454 0.4878 0.4344 0.292 0.2571 0.8868 0.8766
TimesNet 0.489 0.4613 0.8788 0.8248 0.5654 0.5223 0.6151 0.5569 0.4492 0.4091 0.2564 0.2204 0.7233 0.6843
OFA 0.5476 0.5042 0.8462 0.7761 0.6503 0.6112 0.8484 0.8265 0.8294 0.8134 0.3499 0.2869 0.8236 0.8421
FITS 0.6748 0.6674 0.8808 0.8092 0.659 0.715 0.8803 0.8794 0.9316 0.9254 0.4164 0.3766 0.9639 0.9672
Table 6. Performance comparison under point-wise PA (Point Adjustment), reduced-length PA and event-wise PA when using naive training schema.
Method AIOPS NAB TODS WSD Yahoo UCR NEK
$F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$
MatrixProfile 0.1915 0.0281 0.0171 0.7873 0.1769 0.0567 0.5284 0.2058 0.1389 0.1233 0.0371 0.0279 0.3079 0.2051 0.1678 0.7992 0.5066 0.4438 0.6859 0.3163 0.232
SAND 0.271 0.0674 0.0397 0.7007 0.3036 0.1179 0.5372 0.2544 0.1879 0.1761 0.0952 0.0839 0.5718 0.468 0.4235 0.7489 0.5494 0.525 0.5641 0.1758 0.1106
Sub-LOF 0.7273 0.3764 0.2805 0.9787 0.6982 0.6221 0.7997 0.5791 0.4795 0.8683 0.7194 0.6585 0.572 0.4875 0.4453 0.8811 0.5806 0.4917 0.884 0.6807 0.6171
AR 0.8775 0.7517 0.6944 0.9982 0.9035 0.78 0.8312 0.7472 0.6987 0.9624 0.868 0.8225 0.9448 0.9195 0.9026 0.7897 0.4454 0.3693 0.9887 0.9555 0.9249
LSTMAD-$\alpha$ 0.9296 0.8081 0.7671 0.9912 0.8984 0.7852 0.8226 0.7015 0.6582 0.9871 0.9381 0.9009 0.6137 0.5496 0.5244 0.7894 0.4836 0.4043 0.999 0.9971 0.9964
LSTMAD-$\beta$ 0.9288 0.8034 0.763 0.9906 0.9031 0.79 0.8724 0.7589 0.7206 0.9861 0.939 0.9021 0.6013 0.5366 0.5102 0.7713 0.4366 0.3562 0.9865 0.9744 0.9692
AE 0.8778 0.7222 0.6743 0.9921 0.8751 0.7738 0.8402 0.6842 0.6404 0.9809 0.918 0.8851 0.67 0.612 0.5891 0.7074 0.4198 0.3606 0.9938 0.9801 0.9698
EncDecAD 0.9015 0.7583 0.7022 0.99 0.8752 0.77 0.7346 0.6359 0.6156 0.9826 0.9285 0.8945 0.5682 0.5018 0.4767 0.6685 0.3178 0.2541 0.9085 0.755 0.7015
Donut 0.8595 0.6957 0.6299 0.9834 0.8397 0.6938 0.8655 0.7207 0.6826 0.9637 0.8787 0.8326 0.7306 0.6813 0.6598 0.7514 0.4409 0.358 0.9968 0.9885 0.9879
FCVAE 0.9183 0.7851 0.7364 0.9948 0.874 0.7933 0.8658 0.7145 0.6689 0.9689 0.9087 0.8695 0.7334 0.6883 0.6699 0.8357 0.5873 0.5126 0.9733 0.8977 0.8733
SRCNN 0.4069 0.1627 0.1338 0.9079 0.5802 0.4275 0.6298 0.3042 0.2066 0.401 0.1785 0.1221 0.2299 0.1261 0.0849 0.7496 0.3227 0.2209 0.778 0.4365 0.3082
AnomalyTransformer 0.6235 0.3728 0.3124 0.9757 0.7739 0.6117 0.5249 0.2827 0.2123 0.4414 0.2984 0.2462 0.2156 0.145 0.1193 0.7251 0.3104 0.2076 0.683 0.3997 0.3086
TFAD 0.602 0.3551 0.2638 0.8609 0.5746 0.4979 0.6686 0.4526 0.3751 0.8572 0.7561 0.7113 0.8277 0.781 0.7581 0.6456 0.3398 0.2846 0.8653 0.625 0.5245
TranAD 0.7696 0.6486 0.6051 0.9975 0.9101 0.8133 0.5131 0.2531 0.2127 0.7616 0.6738 0.6468 0.6222 0.562 0.539 0.6101 0.2527 0.1946 0.9481 0.9071 0.8924
TimesNet 0.8018 0.6737 0.6379 0.9875 0.8419 0.7488 0.6234 0.3695 0.3048 0.9354 0.8684 0.8358 0.487 0.4444 0.4279 0.621 0.2493 0.1873 0.9807 0.9335 0.9098
OFA 0.881 0.6891 0.615 0.9962 0.8869 0.7841 0.6928 0.6263 0.5811 0.9564 0.8784 0.8344 0.7979 0.7512 0.7315 0.7286 0.4169 0.3489 0.9839 0.9471 0.927
FITS 0.9108 0.7139 0.6324 0.9962 0.8617 0.7447 0.7543 0.5371 0.4873 0.9705 0.8796 0.8377 0.8173 0.7564 0.7244 0.7567 0.4092 0.3284 0.9806 0.8975 0.8462
Table 7. Performance comparison under point-wise PA (Point Adjustment), reduced-length PA and event-wise PA when using all-in-one training schema.
Method AIOPS NAB TODS WSD Yahoo UCR NEK
$F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$
MatrixProfile - - - - - - - - - - - - - - - - - - - - -
SAND - - - - - - - - - - - - - - - - - - - - -
Sub-LOF - - - - - - - - - - - - - - - - - - - - -
AR 0.9141 0.8214 0.7838 0.9978 0.8881 0.7452 0.8696 0.7859 0.745 0.9863 0.9276 0.8883 0.964 0.948 0.9414 0.7983 0.4636 0.3792 0.9897 0.9612 0.9395
LSTMAD-$\alpha$ 0.924 0.8141 0.775 0.9891 0.8492 0.7363 0.8712 0.7538 0.7247 0.9899 0.9467 0.9119 0.8342 0.8 0.7861 0.8915 0.6144 0.5098 0.9795 0.9349 0.9204
LSTMAD-$\beta$ 0.9235 0.8151 0.7744 0.989 0.8484 0.7367 0.8771 0.7495 0.7121 0.99 0.9445 0.9091 0.8455 0.8114 0.7953 0.8691 0.6101 0.5113 0.9906 0.961 0.9451
AE 0.8994 0.7448 0.6908 0.9838 0.8375 0.7336 0.8997 0.7712 0.7129 0.9844 0.9192 0.8835 0.8551 0.8205 0.805 0.7053 0.3644 0.2848 0.9892 0.9609 0.9489
EncDecAD 0.8527 0.7003 0.6405 0.9866 0.8336 0.7321 0.8367 0.6534 0.596 0.9274 0.8353 0.7969 0.6486 0.5842 0.5581 0.7167 0.3333 0.2486 0.9878 0.9534 0.9305
Donut 0.818 0.5827 0.5197 0.988 0.7966 0.5949 0.9006 0.836 0.8111 0.9196 0.7666 0.7004 0.7144 0.6498 0.6209 0.7711 0.4 0.2896 0.8915 0.5801 0.4765
FCVAE 0.8831 0.7593 0.7066 0.9962 0.8857 0.7824 0.8989 0.8526 0.8321 0.9382 0.8121 0.7558 0.8782 0.8537 0.8408 0.8339 0.4766 0.3665 0.9613 0.8148 0.7436
SRCNN 0.8444 0.6672 0.6023 0.9041 0.635 0.5129 0.5706 0.2598 0.1749 0.8906 0.7742 0.7165 0.2591 0.1728 0.1361 0.813 0.3791 0.2668 0.7491 0.6173 0.5637
AnomalyTransformer 0.5024 0.3656 0.2962 0.9784 0.7322 0.5551 0.5168 0.2631 0.1969 0.7379 0.5838 0.494 0.328 0.2377 0.1939 0.6563 0.2372 0.1443 0.8599 0.666 0.5899
TFAD 0.7809 0.4889 0.3698 0.8816 0.4679 0.3274 0.7778 0.5626 0.4776 0.9436 0.8016 0.7128 0.9656 0.9361 0.9176 0.7612 0.4941 0.439 0.8196 0.4103 0.2529
TranAD 0.7401 0.6561 0.614 0.9965 0.8567 0.7032 0.7134 0.5624 0.5413 0.8373 0.7409 0.702 0.582 0.5211 0.4978 0.6075 0.2483 0.1949 0.971 0.9571 0.9546
TimesNet 0.7452 0.4988 0.4227 0.9879 0.8544 0.7505 0.7245 0.5032 0.424 0.7342 0.539 0.4556 0.5644 0.5341 0.5185 0.547 0.1856 0.1185 0.9796 0.9257 0.8947
OFA 0.7934 0.5544 0.4651 0.9879 0.861 0.7053 0.745 0.6187 0.5494 0.926 0.8111 0.7568 0.8251 0.766 0.7367 0.7415 0.2888 0.1914 0.9429 0.7642 0.6954
FITS 0.8807 0.6986 0.623 0.9964 0.8683 0.7508 0.8831 0.7146 0.6548 0.9684 0.8788 0.8315 0.9435 0.9261 0.9184 0.7477 0.428 0.3584 0.9895 0.9625 0.9447
Table 8. Performance comparison under point-wise PA (Point Adjustment), reduced-length PA and event-wise PA when using zero-shot learning schema.
Method AIOPS NAB TODS WSD Yahoo UCR NEK
$F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$ $F1_{p}$ $F1_{r}$ $F1_{e}$
MatrixProfile - - - - - - - - - - - - - - - - - - - - -
SAND - - - - - - - - - - - - - - - - - - - - -
Sub-LOF - - - - - - - - - - - - - - - - - - - - -
AR 0.8991 0.8013 0.7603 0.9981 0.8967 0.7467 0.9192 0.7975 0.7697 0.9886 0.926 0.8765 0.9535 0.9366 0.9305 0.8221 0.487 0.4046 0.9893 0.9641 0.9371
LSTMAD-$\alpha$ 0.9109 0.8184 0.7866 0.9966 0.8593 0.7127 0.8163 0.677 0.6485 0.9902 0.9349 0.8899 0.86 0.8277 0.8147 0.8782 0.5927 0.4898 0.9879 0.9539 0.9478
LSTMAD-$\beta$ 0.9073 0.8073 0.7763 0.9967 0.8556 0.706 0.7793 0.593 0.5435 0.9892 0.9357 0.8888 0.8312 0.7942 0.7785 0.8367 0.5603 0.4681 0.9878 0.9613 0.9573
AE 0.8998 0.7689 0.7251 0.9961 0.8538 0.7133 0.8521 0.7105 0.6829 0.9787 0.8969 0.8603 0.8865 0.8595 0.8472 0.6793 0.3239 0.2497 0.9909 0.9591 0.9451
EncDecAD 0.8873 0.7537 0.7097 0.9952 0.85 0.725 0.8139 0.7089 0.6862 0.9373 0.8306 0.7922 0.6596 0.5958 0.5668 0.7232 0.3209 0.243 0.986 0.9441 0.9332
Donut 0.8766 0.6863 0.6159 0.9928 0.7187 0.4022 0.8324 0.6223 0.5523 0.924 0.7934 0.722 0.7724 0.7034 0.6705 0.7511 0.4148 0.3297 0.9203 0.6204 0.5003
FCVAE 0.9096 0.7983 0.7679 0.9991 0.9385 0.8133 0.8611 0.7987 0.7708 0.9145 0.7707 0.7184 0.8307 0.796 0.7778 0.7691 0.4101 0.325 0.9551 0.7979 0.7439
SRCNN 0.8792 0.6888 0.626 0.9596 0.6421 0.6092 0.5896 0.2977 0.2307 0.8633 0.7414 0.6861 0.2384 0.1437 0.1051 0.8107 0.3582 0.2269 0.7105 0.5261 0.448
AnomalyTransformer 0.5934 0.4361 0.351 0.9526 0.5601 0.4222 0.388 0.2199 0.175 0.533 0.3027 0.2251 0.4807 0.4302 0.4069 0.7598 0.2749 0.1744 0.6492 0.5376 0.4969
TFAD 0.805 0.4659 0.3241 0.9144 0.2806 0.2169 0.8302 0.6352 0.5554 0.9639 0.8902 0.833 0.9586 0.9312 0.9169 0.7316 0.4689 0.4156 0.8352 0.4527 0.3145
TranAD 0.6288 0.4662 0.4031 0.9989 0.9333 0.8 0.5918 0.4013 0.3485 0.7219 0.62 0.591 0.5579 0.4878 0.458 0.6554 0.292 0.2332 0.9258 0.8868 0.8842
TimesNet 0.7622 0.489 0.4329 0.9974 0.8788 0.73 0.7976 0.5654 0.497 0.7777 0.6151 0.5544 0.5038 0.4492 0.4233 0.6406 0.2564 0.1945 0.9467 0.7233 0.5973
OFA 0.7912 0.5476 0.4538 0.9965 0.8462 0.6971 0.8413 0.6503 0.5877 0.9423 0.8484 0.8032 0.8749 0.8294 0.8083 0.7771 0.3499 0.2368 0.9497 0.8236 0.7615
FITS 0.8772 0.6748 0.586 0.9973 0.8808 0.7282 0.8688 0.659 0.593 0.9678 0.8803 0.8406 0.9475 0.9316 0.9233 0.7395 0.4164 0.339 0.9899 0.9639 0.9469
Table 9. Performance differences on all datasets using different learning schemas. Best scores are highlighted in bold, and second-best scores are highlighted in bold and underlined. Performance improvements exceeding 5% are marked with $\uparrow$, while performance declines exceeding 5% are marked with $\downarrow$.
Method AIOPS NAB TODS WSD Yahoo UCR NEK

### v3

@@ V-C Performance under Delay-Constraint
V-D Trade-off Between Performance, Cost, and Efficiency
When deploying algorithms in real-world scenarios, it is often necessary to strike a balance between performance, storage costs, and inference speed. Notably different from other surveys, in this benchmark, we prefer inference time over training time due to the practical needs of anomaly detection. Each factor plays a critical role in determining the feasibility and practicality of the deployed algorithm. The trade-offs between these considerations are important to ensure an efficient and effective deployment that meets the specific requirements and constraints of the application. For example, real-time monitoring or critical systems may prioritize efficiency to detect anomalies promptly and respond quickly, while IoT (Internet of Things) devices with limited on-chip memory have to focus on the performance of models with a small number of parameters. We exclude the statistical methods, as their inference time is not a fixed value (polynomially correlated with the volume of historical data). Also, we exclude TFAD because it does not align with the real-time manner. As shown in Fig. 13, the inference time for a single sample is far less than 50 milliseconds for all methods under our experimental platform. The learning-based AR exhibits the most favorable gain-to-cost ratio among all methods. FCVAE takes the longest time to inference due to its multi-epoch MCMC process. EncDecAD has the second longest inference time because it utilizes LSTM to perform inference for 100 time steps. In addition to providing guidance for practical applications, we aim for this perspective and toolkit to assist in discovering the scaling law in the field of time series anomaly detection under more reasonable inductive bias and larger parameter spaces (RI 14).
Figure 13: The triad of the model’s performance, cost, and efficiency trade-off. The x-axis represents the total inference time (batch size set to 1) of the detectors on AIOPS_7103fa0f-cac4-314f-addc-866190247439 (around 140,000 samples) under default parameters. The y-axis represents the average performance of the models on all datasets under the all-in-one learning schema. The size of the scatter points denotes each method’s cube root of the number of parameters.
VI TimeSeriesBench Toolkit
The overview of the toolkit is shown in Fig. 14. Compared to existing benchmark suites on univariate time series anomaly detection [14, 13, 15, 54], our framework provides more possibilities for diverse experiment setups. As shown in Table III, in addition to including rich datasets and diverse learning schemas, the toolkit considers the need for more kinds of evaluation criteria. We provide more famous evaluation criteria as built-in criteria. Thus researchers can conveniently compare the strengths and weaknesses of different criteria.
Despite the built-in settings, TimeSeriesBench toolkit also takes the requirements for extensibility into consideration. We provide flexible interfaces for dataset, method, evaluation criteria, runtime statistic (RT), and learning schema, while existing benchmarks only provides only parts of these interfaces. The dataset interface and algorithm interfaces are provided to allow evaluations on new or private datasets and novel methods under all learning schemas. All plots of raw curves and scores are saved for specific investigation. Previously mentioned flaws in evaluation criteria have sparked a research fervor and several latest works are dedicated to providing evaluation criteria according to different assumptions [32, 55, 35, 34, 30, 36]. Thus we also expose an interface for swiftly developing neo-criteria for evaluation grounded in realistic assumptions and evaluating the performance of all methods w.r.t various datasets. In some scenarios, there might be a trade-off between model accuracy and model training/testing cost. Therefore, we also provide a runtime statistics interface to conveniently track and analyze the runtime information of the model. Also, you can design your own training, test and other dataflows via implementing learning schema interface.
Figure 14: Overview of TimeSeriesBench Toolkit.
Owing to the flexible framework, our benchmark suite can incorporate real-world scenarios and challenges, aiming to replicate the complexities and nuances encountered in practical applications. This enables researchers and practitioners to gain deeper insights into the strengths and weaknesses of various deep learning approaches, and is well-prepared for the emergence and development of Foundation Models (FMs) in the field of time series anomaly detection. In the toolkit, we provide four kinds of workflows to meet different research needs for time series anomaly detection: algorithms benchmarking, algorithm development, evaluation criteria development, and performance analysis. For more information about TimeSeriesBench, please refer to our repository and the website of the leaderboard.
TABLE III: Comparison among TimeSeriesBench and existing UTS anomaly detection benchmark suites. TimeSeriesBench presents more angles for performance evaluation, meanwhile offering user-friendly interfaces for up-to-date model and evaluation criterion developments.
Suites Exathlon[14] TODS[13] TSB[15] TimeEval[54] Ours
Data Source Real-World
Synthetic
Learning Schema One-by-one
All-in-one
Zero-shot
Supported Eval criteria Point-based [^4] Point-based criteria treat each individual data point as a separate sample, disregarding the holistic characteristics of the anomaly segment.
Range-based [^5] Range-based criteria incorporate segment-level features, for instance, the detection latency, into the evaluation.
Event-based
Extensibility Dataset
Method
Eval criteria
RT Statistics
Learning Schema
VII Threats to validity
As a benchmark, our research is susceptible to several common threats that can compromise the validity and reliability of our findings, including dataset, algorithm settings and the applicable scenarios.
Dataset. Although we have adopted six well-known datasets and have released a real-world dataset, the scenarios covered by these datasets are still limited, so the evaluation results in other scenarios may vary from the main results. We will supplement more datasets in the future to make the evaluation results more generalizable.
Algorithms settings. Due to time and computational resource constraints, we evaluated the algorithms using their default parameter settings in the experiments. If the hyperparameters are fine-tuned for each scenario, some methods might achieve better performance.
Applicable scenarios. In this paper we focus on real-time anomaly detection scenarios, and introduce a more robust criterion for evaluation based on real-time detection demands. All results and analysis are based on this setting. However, for offline detection tasks that do not have real-time requirements and require precise detection of the duration of anomalies, other evaluation criteria such as VUS [34] may be more suitable. You can specify this criterion as the evaluation criteria in TimeSeriesBench toolkit to obtain results that are more tailored to your specific application scenario.
VIII Conclusion AND FUTURE DIRECTIONS
In this paper, we propose TimeSeriesBench, a comprehensive and application-oriented benchmark for evaluating the performance of existing and emerging UTS anomaly detection methods. TimeSeriesBench takes into account existing industrial concerns and conducts a comprehensive performance evaluation of some well-known and latest methods under settings that meet industrial requirements. Also, it offers unprecedented perspectives for measuring algorithm performance, meanwhile laying a solid foundation for the development of Foundation Models in the field. Moreover, TimeSeriesBench includes a user-friendly toolkit and leaderboard, offering easy-to-use interfaces that allow researchers and practitioners to focus on advancing their algorithms without getting entangled in repetitive tasks. We intend to incorporate more latest methods/evaluation criteria and employ more high-quality data. Also, we will continue to monitor the latest developments in general foundation models for time series within the realm of anomaly detection, adapting them accordingly within our benchmark. Additionally, whether you propose deep learning or statistical methods, we welcome your participation in our leaderboard to help drive the development of the time series anomaly detection community.
ACKNOWLEDGMENTS
This work was supported by the National Key Research and Development Program of China (No. 2021YFE0111500), in part by the Chinese Academy of Sciences (No. 241711KYSB20200023), in part by the National Natural Sci- ence Foundation of China (No. 62202445), and in part by the State Key Program of the National Natural Science Foundation of China under Grant 62321166652.